OpenStudy (anonymous):

Using the sieve of Eratosthenes, circle the number two, now cross out every multiple of two and write an algebraic expression for the numbers u crossed out

OpenStudy (agreene):

there isnt really a question here, but what you've written does seem to be the seive of Erato.. if I remember it correctly. just follow the directions and im sure you'll be fine.

OpenStudy (agreene):

i suppose, it is asking a question, to write an expression for all the even numbers (as that is the 1st set of crossed out numbers, and all multiples of 2 are all the evens) n=2k would be that relation. ex: 2*1=2 2*2=4 2*3=6 ... 2*(n-1)=2n-2 2n=2n